### v1

@@ Abstract
The EM (Expectation-Maximization) algorithm is regarded as an MM (Majorization-Minimization) algorithm for maximum likelihood estimation of statistical models. Expanding this view, this paper demonstrates that by choosing an appropriate probability distribution, even nonstatistical optimization problem can be cast as a negative log-likelihood-like minimization problem, which can be approached by an EM (or MM) algorithm. When a polynomial objective is optimized over a simple polyhedral feasible set and an exponential family distribution is employed, the EM algorithm can be reduced to a natural gradient descent of the employed distribution with a constant step size. This is demonstrated through three examples. In this paper, we demonstrate the global convergence of specific cases with some exponential family distributions in a general form. In instances when the feasible set is not sufficiently simple, the use of MM algorithms can nevertheless be adequately described. When the objective is to minimize a convex quadratic function and the constraints are polyhedral, global convergence can also be established based on the existing results for an entropy-like proximal point algorithm.
keywords:
@@
EM algorithm, MM algorithm, polynomial objective, natural gradient descent, exponential family distributions
1 Introduction
@@
After being rediscovered many times by predecessors, the EM (Expectation-Maximization) algorithm was coined and popularized by Dempster, Laird, and Rubin [10] in late 1970s as a framework used in estimating statistical models, in particular, when those depend on latent random variables. See, for example, [25] for plenty of examples of successful applications of the framework and extensive references therein. In the context of mathematical optimization, the EM algorithm can be viewed as a special example of MM (Majorization-Minimization) algorithm (e.g., [15, 22]) or a proximal-like algorithm that employs the Kullback-Leibler divergence (e.g., [32]). Focusing on such an MM aspect, we study solving a general, i.e., not necessarily statistical, optimization problem via an EM algorithm (without observed data) by rewriting it to another optimization of log-likelihood-like function with the help of a certain probability distribution and by reading the parameters of the distribution as the decision variables of optimization problems in hand.
@@
Table 1 summarizes examples of such combination of optimization problem and exponential family distribution explored in this article. In particular, for the first three cases (i), (ii), and (iii) in Table 1, we show that the EM algorithm results in a gradient descent algorithm with a natural gradient [1] with a constant step size. Even when the constraints are not simple, as (iv) to (vi), the EM scheme yields MM algorithms which can further derive EM gradient algorithms [21], under which a bit more practical presentations of interior point methods can be derived.
@@
Table 1: Problem types and used probability distributions for EM algorithms
@@
(iii) polynomial simplex polynomial yes
@@
(vi) quadratic polyhedron Poisson$+$normal (const.cov.) no
@@
The article [5] concentrates on the EM algorithms with no missing data, which is usually presumed when dealing with the EM algorithms. It asserts that the MM aspect is indispensable for the EM algorithms as optimization algorithms. This paper makes a similar claim, but in contrast to [5], it focuses on the EM algorithms with no observed data, i.e., all data is missing. This contrast stems from the difference in stand points that [5] is aimed at statisticians, while this paper is aimed at a broader readership including optimization researchers.
@@
The next section describes basic strategy and required components of the proposed scheme. Section 3 shows properties of the scheme for the case where the involved distribution is a member of the exponential family. Section 4 presents three examples (i) to (iii) in Table 1. Section 5 is devoted to further examples (iv) and (vi), which are based on shifted exponential family distributions. Section 6 concludes the paper. The majority of the proofs of propositions are located in the appendix, with the intention of enhancing the readability of the manuscript. The example (v) is also described in the appendix.
@@
2 EM scheme for general optimization problem
$\underset{\bm{\theta}\in\Theta}{\min}~{}$ $f(\bm{\theta})$ (1)
$u(\bm{\theta}|\bar{\bm{\theta}})$ $\geq f(\bm{\theta}),\quad\mbox{for any }\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\Theta,$ (2)
$u(\bar{\bm{\theta}}|\bar{\bm{\theta}})$ $=f(\bar{\bm{\theta}}),\quad\mbox{for any }\bar{\bm{\theta}}\in\mathrm{int}\Theta.$ (3)
Let $\bm{X}$ be a vector of random variables that has a probability density $p_{\bm{x}}$ such that $\mathbb{P}\{\bm{X}\leq\bm{x}|\bm{\theta}\}=\int_{\bm{x}\in\Xi}p_{\bm{x}}(\bm{\theta})\mathrm{d}\bm{x}$, a probability mass $p_{\bm{x}}(\bm{\theta})\equiv\mathbb{P}\{\bm{X}=\bm{x}|\bm{\theta}\}$, or their hybrid, where $\bm{\theta}\in\Theta\subset\mathbb{R}^{p}$ is the vector of parameters of the probability distribution and $\Xi$ is the support of $\bm{X}$. Assume that for (1) and $p_{\bm{x}}$ satisfying $\mathrm{int}\Theta=\{\bm{\theta}\in\mathbb{R}^{p}|p_{\bm{x}}(\bm{\theta})>0\}$, there is a positive-valued function $G:\Xi\to(0,\infty)$ satisfying
$f(\bm{\theta})=-\ln\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big{[}G(\bm{X})\big{]}\quad\mbox{for all }\bm{\theta}\in\mathrm{int}\Theta,$ (4)
where $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big{[}\cdot\big{]}$ denotes the expectation under the distribution $p_{\bm{x}}(\bm{\theta})$. (For the expectation computation in this article, we only describe the case where $p_{\bm{x}}$ is a probability density because it is straightforward to get results for the other cases, for example, by replacing the integral with a discrete summation when $p_{\bm{x}}$ is a probability mass.) While we consider minimizing the negative log-likelihood-like function (4) for (1), it corresponds to the EM algorithm where no data point is observed and only the distribution of latent variables is assumed, which is in contrast to the ordinary context or [5]. As will be shown below, however, this set-up is enough to link the EM algorithm to (1). Some examples of the tuple $(f,\Theta,p_{\bm{x}},G)$ will be presented in Sections 4 and 5.
For a tuple $(f,\Theta,p_{\bm{x}},G)$ satisfying (4), we see that
$u(\bm{\theta}|\bar{\bm{\theta}})$ $:=-\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\ln\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta})}{q_{\bm{x}}(\bar{\bm{\theta}})}\mathrm{d}\bm{x}$ (5)
satisfies (2) since, for another density $q_{\bm{x}}$ (equivalent to $p_{\bm{x}}$), we have for any $\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\Theta$,
$f(\bm{\theta})$ $=-\ln\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta})}{q_{\bm{x}}(\bar{\bm{\theta}})}\mathrm{d}\bm{x}\leq-\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\ln\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta})}{q_{\bm{x}}(\bar{\bm{\theta}})}\mathrm{d}\bm{x}\equiv u(\bm{\theta}|\bar{\bm{\theta}}),$ (6)
where the inequality holds due to the Jensen’s inequality. Letting $s(\bm{x}|\bm{\theta}):=G(\bm{x})p_{\bm{x}}(\bm{\theta}),$ we see that the nonnegative gap, denoted $h(\bm{\theta}|\bar{\bm{\theta}})$, between $u(\bm{\theta}|\bar{\bm{\theta}})$ and $f(\bm{\theta})$ equals the Kullback-Leibler (KL) divergence $\mathrm{KL}\big{(}\bm{q}\big{\|}\bm{p}\big{)}:=\int_{\bm{x}\in\Xi}q_{\bm{x}}\ln\frac{q_{\bm{x}}}{p_{\bm{x}}}\mathrm{d}\bm{x}$ of the distribution $\bm{q}:=\big{(}q_{\bm{x}}(\bar{\bm{\theta}})\big{)}_{\bm{x}\in\Xi}$ relative to $\bm{p}:=\big{(}\frac{s(\bm{x}|\bm{\theta})}{\int_{\bm{\xi}}s(\bm{\xi}|\bm{\theta})\mathrm{d}\bm{\xi}}\big{)}_{\bm{x}\in\Xi}$, namely,
$h(\bm{\theta}|\bar{\bm{\theta}}):=u(\bm{\theta}|\bar{\bm{\theta}})-f(\bm{\theta})$ $=\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\ln\frac{q_{\bm{x}}(\bar{\bm{\theta}})}{\left(\displaystyle\frac{s(\bm{x}|\bm{\theta})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta})\mathrm{d}\bm{x}}\right)}\mathrm{d}\bm{x}\geq 0.$ (7)
Therefore, if we set $\bar{\bm{\theta}}$ so that $q_{\bm{x}}(\bar{\bm{\theta}})=\frac{s(\bm{x}|\bar{\bm{\theta}})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bar{\bm{\theta}})\mathrm{d}\bm{x}}$ is fulfilled, the function $u$ satisfies (3).
@@ 2 EM scheme for general optimization problem
$q_{\bm{x}}(\bm{\theta}^{(t-1)})$ $\leftarrow\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}\bm{x}}\equiv\frac{G(\bm{x})}{\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t-1)})}[G(\bm{X})]}p_{\bm{x}}(\bm{\theta}^{(t-1)}),\qquad\mbox{for }\bm{x}\in\Xi,$ (8)
which minimizes the KL divergence relative to $\big{(}\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}\bm{x}}\big{)}_{\bm{x}}$, ensuring $u(\bm{\theta}^{(t-1)}|\bm{\theta}^{(t-1)})=f(\bm{\theta}^{(t-1)}).$ The surrogate function $u(\bm{\theta}|\bm{\theta}^{(t-1)})$ is then minimized:
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)}).$ (9)
The EM algorithm repeats the two steps as in Algorithm 1. If the minimizer (9) is found, it is a solution whose objective value is no worse than the incumbent since we have
@@
Algorithm 1 Prototype of EM algorithm for optimization problem (1)
1:Let $(f,\Theta,p_{\bm{x}},G)$ be a tuple satisfying (4), and $\bm{\theta}^{(0)}\in\mathrm{int}\Theta$, and set $t\leftarrow 1$.
@@
3:     $q_{\bm{x}}(\bm{\theta}^{(t-1)})\leftarrow\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}\bm{\xi}}\equiv\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}G(\bm{\xi})p_{\bm{\xi}}(\bm{\theta}^{(t-1)})\mathrm{d}\bm{\xi}},\qquad\mbox{for }\bm{x}\in\Xi$
4:     $\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)})\equiv\arg\min_{\bm{\theta}\in\Theta}\Big{\{}-\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bm{\theta}^{(t-1)})\ln s(\bm{x}|\bm{\theta})\mathrm{d}\bm{x}\Big{\}}.$
@@
$\underset{\bm{\theta}\in\Theta}{\min}~{}$ $\Phi(\bm{\theta}):=f(\bm{\theta})+\rho(\bm{\theta}),$ (10)
where $\rho$ is a function on $\Theta$. For any surrogate function $u(\cdot|\cdot)$ for $f(\cdot)$, $U(\bm{\theta}|\bar{\bm{\theta}}):=u(\bm{\theta}|\bar{\bm{\theta}})+\rho(\bm{\theta})$ is a surrogate function for $\Phi(\bm{\theta})$ and an MM algorithm for (10) can be constructed by simply replacing the M-step (line 4) of Algorithm 1 with $\bm{\theta}^{(t)}\leftarrow\arg\min\limits_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)})+\rho(\bm{\theta})$. An example of this extension will be given in Section 4.1.1.
3 EM scheme with exponential family distributions

### v2

@@ Abstract.
The EM (Expectation–Maximization) algorithm is regarded as an MM (Majorization–Minimization) algorithm for maximum likelihood estimation of statistical models. Expanding this view, this paper demonstrates that by choosing an appropriate probability distribution, even nonstatistical optimization problem can be cast as a negative log-likelihood-like minimization problem, which can be approached by an EM (or MM) algorithm. When a polynomial objective is optimized over a simple polyhedral feasible set and an exponential family distribution is employed, the EM algorithm can be reduced to a natural gradient descent of the employed distribution with a constant step size. This is demonstrated through three examples. In this paper, we demonstrate the global convergence of specific cases with some exponential family distributions in a general form. In instances when the feasible set is not sufficiently simple, the use of MM algorithms can nevertheless be adequately described. When the objective is to minimize a convex quadratic function and the constraints are polyhedral, global convergence can also be established based on the existing results for an entropy-like proximal point algorithm.
Disclaimer: The first author of this paper hereby states that his affiliated institution has had no direct involvement in the research and bears no responsibility for the findings or conclusions of this study.
@@
Keywords: EM algorithm, MM algorithm, polynomial objective, natural gradient descent, exponential family distributions
1. Introduction
@@
After being rediscovered many times by predecessors, the EM (Expectation–Maximization) algorithm was coined and popularized by Dempster, Laird, and Rubin [13] in late 1970s as a framework used in estimating statistical models, in particular, when those depend on latent random variables. See, for example, [28] for plenty of examples of successful applications of the framework and extensive references therein. In the context of mathematical optimization, the EM algorithm can be viewed as a special example of MM (Majorization–Minimization) algorithm (e.g., [18, 25]) or a proximal-like algorithm that employs the Kullback–Leibler divergence (e.g., [36]). Focusing on such an MM aspect, we study solving a general, i.e., not necessarily statistical, optimization problem via an EM algorithm (without observed data) by rewriting it to another optimization of log-likelihood-like function with the help of a certain probability distribution and by reading the parameters of the distribution as the decision variables of optimization problems in hand.
@@
Table 1.1 summarizes examples of such combination of optimization problem and exponential family distribution explored in this article. In particular, for the first three cases (i), (ii), and (iii) in Table 1.1, we show that the EM algorithm results in a gradient descent algorithm with a natural gradient [4] with a constant step size. Even when the constraints are not simple, as (iv) to (vi), the EM scheme yields MM algorithms which can further derive EM gradient algorithms [24], under which a bit more practical presentations of interior point methods can be derived.
@@
Table 1.1. Problem types and used probability distributions for EM algorithms
@@
(iii) polynomial simplex multinomial yes
@@
(vi) quadratic polyhedron Poisson++normal (const.cov.) no
@@
Researchers have employed the optimization of parameters of probability mass functions belonging to the exponential family in contexts other than purely statistical estimation, such as maximum likelihood estimation. For example, [32] presented a methodology for automatically estimating the structure of a neural network by minimizing the expected loss under a multivariate Bernoulli distribution, where the optimization variables are the parameters of the distribution, using a stochastic gradient descent method based on the natural gradient. In contrast to those studies, this paper does not specify an application. Rather, it systematically demonstrates that by selecting an appropriate exponential distribution family, it is possible to construct an EM (MM) algorithm and a natural gradient descent method, even for a general class of optimization problems that do not contain any statistical elements.
In terms of focusing on the MM aspect of the EM algorithm and extending it to general optimization problems, the paper by Becker et al. [7] is closer to ours, even though their argument starts from the statistical EM algorithm. To make a similar claim, they present the EM algorithms in the absence of missing data, which is typically assumed when working with the EM algorithms. In contrast to [7] and other statistical EM algorithms, this paper focuses on the EM algorithms with no observed data, i.e., all data is missing, to extend to non-statistical optimization problems. This contrast stems from the difference in stand points that [7] is aimed at statisticians, while this paper is aimed at a broader readership including optimization researchers. In fact, the present paper puts forth a novel perspective in which an optimization problem devoid of statistical elements can be artificially linked to log-likelihood-like optimization, and through the EM algorithm for that, we derive an MM algorithm for the original optimization problem.
@@
The next section describes basic strategy and required components of the proposed scheme. Section 3 shows properties of the scheme for the case where the involved distribution is a member of the exponential family. Section 4 presents three examples (i) to (iii) in Table 1.1. Section 5 is devoted to further examples (iv) and (vi), which are based on shifted exponential family distributions. Section 6 concludes the paper. The majority of the proofs of propositions are located in the appendix, with the intention of enhancing the readability of the manuscript. The example (v) is also described in the appendix.
@@
We use small bold letters for indicating vectors and capital bold letters for indicating matrices. In particular, $\bm{0},\bm{1}$ denote the all-zero or all-one vectors, respectively. Small italic letters are used to stand for components of vectors and matrices, e.g., $\bm{\theta}=(\theta_{1},...,\theta_{p})^{\top}$. $\mathbb{R}^{p}_{+}$ and $\mathbb{R}^{p}_{++}$ denote the spaces of nonnegative and positive vectors, respectively. We denote by $\mathbb{S}^{p}$ the space of $p\times p$ real symmetric matrices, by $\mathbb{S}^{p}_{+},\mathbb{S}^{p}_{++}$ the spaces of positive semidefinite and positive definite matrices in $\mathbb{S}^{p}$, respectively. For $\bm{A},\bm{B}\in\mathbb{S}^{p}$, $\bm{A}\succ\bm{B}$ stands for $\bm{A}-\bm{B}\in\mathbb{S}^{p}_{++}$. By $\|\cdot\|$ we denote the Euclidean norm. For $\bm{Q}\in\mathbb{S}^{p}_{++}$, we denote the ellipsoidal norm by $\|\cdot\|_{\bm{Q}}$, i.e., $\|\bm{z}\|_{\bm{Q}}:=\sqrt{\bm{z}^{\top}\bm{Q}\bm{z}}$. We use the abbreviation $p]:=\{1,...,p\}$.
2. EM scheme for general optimization problem
$\underset{\bm{\theta}\in\Theta}{\mbox{minimize}}$~ $f(\bm{\theta})$ (2.1)
$u(\bm{\theta}|\bar{\bm{\theta}})$ $\geq f(\bm{\theta}),\quad\mbox{for any }\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\Theta,$ (2.2)
$u(\bar{\bm{\theta}}|\bar{\bm{\theta}})$ $=f(\bar{\bm{\theta}}),\quad\mbox{for any }\bar{\bm{\theta}}\in\mathrm{int}\Theta.$ (2.3)
Let $\bm{X}$ be a vector of random variables that has a probability density $p_{\bm{x}}$ such that $\mathbb{P}\{\bm{X}\leq\bm{x}|\bm{\theta}\}=\int_{\bm{\xi}\leq\bm{x}}p_{\bm{\xi}}(\bm{\theta})\mathrm{d}\bm{\xi}$, a probability mass $p_{\bm{x}}(\bm{\theta})=\mathbb{P}\{\bm{X}=\bm{x}|\bm{\theta}\}$, or their hybrid, where $\bm{\theta}\in\Theta\subset\mathbb{R}^{p}$ is the vector of parameters of the probability distribution and $\Xi$ is the support of $\bm{X}$. Assume that for (2.1) and $p_{\bm{x}}$ satisfying $\mathrm{int}\Theta=\{\bm{\theta}\in\mathbb{R}^{p}|p_{\bm{x}}(\bm{\theta})>0\}$, there is a positive-valued function $G:\Xi\to(0,\infty)$ satisfying
$f(\bm{\theta})=-\ln\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big[G(\bm{X})\big]\quad\mbox{for all }\bm{\theta}\in\mathrm{int}\Theta,$ (2.4)
where $\mathbb{E}_{p_{\bm{x}}(\bm{\theta})}\big[\cdot\big]$ denotes the expectation under the distribution $p_{\bm{x}}(\bm{\theta})$. (For the expectation computation in this article, we only describe the case where $p_{\bm{x}}$ is a probability density because it is straightforward to get results for the other cases, for example, by replacing the integral with a discrete summation when $p_{\bm{x}}$ is a probability mass.) While we consider minimizing the negative log-likelihood-like function (2.4) for (2.1), it corresponds to the EM algorithm where no data point is observed and only the distribution of latent variables is assumed, which is in contrast to the ordinary context or [7]. As will be shown below, however, this set-up is enough to link the EM algorithm to (2.1). Some examples of the tuple $(f,\Theta,p_{\bm{x}},G)$ will be presented in Sections 4 and 5.
For a tuple $(f,\Theta,p_{\bm{x}},G)$ satisfying (2.4), we see that
$u(\bm{\theta}|\bar{\bm{\theta}})$ $:=-\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\ln\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta})}{q_{\bm{x}}(\bar{\bm{\theta}})}\mathrm{d}\bm{x}$ (2.5)
satisfies (2.2) since, for another density $q_{\bm{x}}$ (equivalent to $p_{\bm{x}}$), we have for any $\bm{\theta},\bar{\bm{\theta}}\in\mathrm{int}\Theta$,
@@ 2. EM scheme for general optimization problem
where the inequality holds due to the Jensen’s inequality. Letting $s(\bm{x}|\bm{\theta}):=G(\bm{x})p_{\bm{x}}(\bm{\theta}),$ we see that the nonnegative gap, denoted $h(\bm{\theta}|\bar{\bm{\theta}})$, between $u(\bm{\theta}|\bar{\bm{\theta}})$ and $f(\bm{\theta})$ equals the Kullback–Leibler (KL) divergence $\mathrm{KL}\big(\bm{q}\big\|\bm{p}\big):=\int_{\bm{x}\in\Xi}q_{\bm{x}}\ln\frac{q_{\bm{x}}}{p_{\bm{x}}}\mathrm{d}\bm{x}$ of the distribution $q_{\bm{x}}=q_{\bm{x}}(\bar{\bm{\theta}}),\bm{x}\in\Xi$, relative to $p_{\bm{x}}=\frac{s(\bm{x}|\bm{\theta})}{\int_{\bm{\xi}}s(\bm{\xi}|\bm{\theta})\mathrm{d}\bm{\xi}},\bm{x}\in\Xi$, that is,
$h(\bm{\theta}|\bar{\bm{\theta}}):=u(\bm{\theta}|\bar{\bm{\theta}})-f(\bm{\theta})$ $=\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bar{\bm{\theta}})\ln\frac{q_{\bm{x}}(\bar{\bm{\theta}})}{\left(\displaystyle\frac{s(\bm{x}|\bm{\theta})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta})\mathrm{d}\bm{\xi}}\right)}\mathrm{d}\bm{x}\geq 0.$ (2.7)
Therefore, if we set $\bar{\bm{\theta}}$ so that $q_{\bm{x}}(\bar{\bm{\theta}})=\frac{s(\bm{x}|\bar{\bm{\theta}})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bar{\bm{\theta}})\mathrm{d}{\bm{\xi}}}$ is fulfilled, the function $u$ satisfies (2.3).
@@
$q_{\bm{x}}(\bm{\theta}^{(t-1)})$ $\leftarrow\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}{\bm{\xi}}}=\frac{G(\bm{x})}{\mathbb{E}_{p_{\bm{x}}(\bm{\theta}^{(t-1)})}[G(\bm{X})]}p_{\bm{x}}(\bm{\theta}^{(t-1)}),\qquad\mbox{for }\bm{x}\in\Xi,$ (2.8)
which minimizes the KL divergence relative to $\big(\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}{\bm{\xi}}}\big)_{\bm{x}}$, ensuring $u(\bm{\theta}^{(t-1)}|\bm{\theta}^{(t-1)})=f(\bm{\theta}^{(t-1)}).$ The surrogate function $u(\bm{\theta}|\bm{\theta}^{(t-1)})$ is then minimized:
$\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)}).$ (2.9)
The EM algorithm repeats the two steps as in Algorithm 1. If the minimizer (2.9) is found, it is a solution whose objective value is no worse than the incumbent since we have
@@
Algorithm 1 Prototype of EM algorithm for optimization problem (2.1)
1:Let $(f,\Theta,p_{\bm{x}},G)$ be a tuple satisfying (2.4), and $\bm{\theta}^{(0)}\in\mathrm{int}\Theta$, and set $t\leftarrow 1$.
@@
3:  $q_{\bm{x}}(\bm{\theta}^{(t-1)})\leftarrow\frac{s(\bm{x}|\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}s(\bm{\xi}|\bm{\theta}^{(t-1)})\mathrm{d}\bm{\xi}}=\frac{G(\bm{x})p_{\bm{x}}(\bm{\theta}^{(t-1)})}{\int_{\bm{\xi}\in\Xi}G(\bm{\xi})p_{\bm{\xi}}(\bm{\theta}^{(t-1)})\mathrm{d}\bm{\xi}},\qquad\mbox{for }\bm{x}\in\Xi$
4:  $\bm{\theta}^{(t)}\leftarrow\arg\min_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)})=\arg\min_{\bm{\theta}\in\Theta}\Big\{-\int_{\bm{x}\in\Xi}q_{\bm{x}}(\bm{\theta}^{(t-1)})\ln s(\bm{x}|\bm{\theta})\mathrm{d}\bm{x}\Big\}.$
@@
$\underset{\bm{\theta}\in\Theta}{\mbox{{minimize}}}$~ $\Phi(\bm{\theta}):=f(\bm{\theta})+\rho(\bm{\theta}),$ (2.10)
where $\rho$ is a function on $\Theta$. For any surrogate function $u(\cdot|\cdot)$ for $f(\cdot)$, $U(\bm{\theta}|\bar{\bm{\theta}}):=u(\bm{\theta}|\bar{\bm{\theta}})+\rho(\bm{\theta})$ is a surrogate function for $\Phi(\bm{\theta})$ and an MM algorithm for (2.10) can be constructed by simply replacing the M-step (line 4) of Algorithm 1 with $\bm{\theta}^{(t)}\leftarrow\arg\min\limits_{\bm{\theta}\in\Theta}u(\bm{\theta}|\bm{\theta}^{(t-1)})+\rho(\bm{\theta})$. An example of this extension will be given in Section 4.1.1.
3. EM scheme with exponential family distributions